OpenStudy (anonymous):

How many units apart are the points (8, 11) and (-1, -1)?

Parth (parthkohli):

 \( \color{Black}{\Rightarrow 9^2 + 12^2 = x^2 }\) Solve for \(x\).

Parth (parthkohli):

This has a distance formula too which is derived from what I said.

OpenStudy (anonymous):

im so confused..

Parth (parthkohli):

I made a right triangle and have marked the two legs. You just have to determine the hypotenuse :)

Parth (parthkohli):

\( \color{Black}{\Rightarrow 81 + 144 = x^2 }\) \( \color{Black}{\Rightarrow 225 = x^2 }\) \( \color{Black}{\Rightarrow x = \sqrt{225} =... }\)

Parth (parthkohli):

What is the square root of 225?

OpenStudy (anonymous):

15?

Parth (parthkohli):

Yes :) that is the distance :D

OpenStudy (anonymous):

woooooo thanks :)

Parth (parthkohli):

Yw :)